OpenStudy (anonymous):

Find a so that x^2 + ax + 1= 0 and x^2 + x + a = 0 have a common root.

OpenStudy (akshay_budhkar):

??? your question?

OpenStudy (amistre64):

a common root being that they cross at a given x=0 point

OpenStudy (akshay_budhkar):

x=1

OpenStudy (akshay_budhkar):

just substitute and see

OpenStudy (amistre64):

find a not x :)

OpenStudy (akshay_budhkar):

its the common root

OpenStudy (anonymous):

Common eg roots alpha and beta, one of them same in both equations.

OpenStudy (akshay_budhkar):

but for all a x=1 is common?

OpenStudy (anonymous):

U r right, I made a booboo...

OpenStudy (anonymous):

x^2 + ax - 1= 0

OpenStudy (akshay_budhkar):

ohhhhh

OpenStudy (akshay_budhkar):

GN guys gotta go! this site is realllllllllly addictive

OpenStudy (amistre64):

x^2 + nx - 1= 0 and x = \(\cfrac{-n \pm\sqrt{n^2+4}}{2}\) and factorize it in x^2 + x + n = 0 x = \(\cfrac{-1\pm\sqrt{1-4n}}{2}\) would this be useful?

OpenStudy (amistre64):

\[-n + \sqrt{n^2+4}=-1+\sqrt{1-4n}\] and fight with it tooth and nail?

OpenStudy (amistre64):

i get n=0 perhaps?

OpenStudy (anonymous):

Doncha just love algebra?

OpenStudy (amistre64):

my link dint linkate

OpenStudy (amistre64):

when a=0; they root at -1

OpenStudy (amistre64):

always remember, when anyone gives you a seemingly difficult question, the answer is either 1 or 0 :)

OpenStudy (anonymous):

???

OpenStudy (amistre64):

x^2 + 0x - 1 and x^2 +x +0 have the same root of x=-1

OpenStudy (anonymous):

U r right, I should have done it ur way...

OpenStudy (anonymous):

Now heres the deal for common roots Lets say two equations \(a_1x^2 +b_1x+c_1=0\) and \(a_2x^2 +b_2x +c_2=0\). \[(b_1c_2 -b_2c_1)(a_1b_2-a_2b_1)=(c_1a_2 - c_2a_1)^2\] So our equations are. \[x^2 +ax+1=0\]\[x^2+x+a=0\] That gives, \[(a^2 - 1)(1-a) = (1-a)^2\] \[a^2 -1 = 1-a \] \[a^2 +a-2 =0\] \[a = \frac{-1 \pm \sqrt{1 + 8}}{2}\]

OpenStudy (amistre64):

my calc 1 teacher said, whenever you have a seemingly impossible equation, the answer is either 1 or 0

OpenStudy (anonymous):

I got -i sqrt 3

OpenStudy (anonymous):

Back later....

hero (hero):

Myininaya, are you busy?

myininaya (myininaya):

sort of what do you want?

hero (hero):

I have a question about your alternative method of grouping

myininaya (myininaya):

one sec...

myininaya (myininaya):

whats your question?

hero (hero):

If you solve 3x^2 +8x + 2 = 0 using your method what do you get?

hero (hero):

That's what I got. Okay. I was just checking

myininaya (myininaya):

lol ok

hero (hero):

Thank you :o)

hero (hero):

I got: \[x = -\frac{4+\sqrt{10}}{3}, \frac{-4+\sqrt{10}}{3}\]

myininaya (myininaya):

thats good :)

myininaya (myininaya):

estudier i like your question i wish i had gotten to it first instead of amistre :(